# Kirby diagrams, trisections and gems of PL 4-manifolds: relationships, results and open problems

### v2

@@ 2 Preliminaries about the involved representations
Obviously ${\partial{\mathbb{Y}}}_{m}^{(\sim)}=\#_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ and, whenever orientability is assumed, the $n$-sphere (resp. the $4$-disk) will be included in the notation #$\#_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{n-1})=\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{n-1})$ (resp. ${\mathbb{Y}}_{m}^{(\sim)}={\mathbb{Y}}_{m}$) as the case $m=0.$
@@ 2.1 Kirby diagrams of orientable $4$-manifolds
In fact, since the existence of exactly one 00-handle can always be assumed, then 1- and 2-handles can be visualized, by positioning on the boundary of the 00-handle, as follows:
@@
In [32], Kirby diagrams representing closed orientable $4$-manifolds (i.e. representing compact $4$-manifolds with boundary homeomorphic to #$\#_{r}(\mathbb{S}^{1}\times\mathbb{S}^{2}),$ with $r\geq 0$) are also called (4-dimensional) Heegaard diagrams.
@@ 2.2 Trisections of closed $4$-manifolds
A $(g;k_{0};k_{1};k_{2})$-trisection diagram is a 4-tuple $(\Sigma;\alpha,\beta,\gamma)$ such that the triples $(\Sigma;\alpha,\beta)$, $(\Sigma;\alpha,\gamma)$ and $(\Sigma;\beta,\gamma)$ are genus $g$ Heegaard diagrams for #$\#_{k_{0}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$, #$\#_{k_{1}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ and #$\#_{k_{2}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ respectively. [^2] Obviously, all $3$- and $4$-dimensional manifolds involved in Definition 2, as well as in the procedure to reconstruct the represented closed $4$-manifold, are orientable or not according to $\Sigma$.
@@
    fill the three boundary components (isomorphic to #$\#_{k_{i}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$, for $i=0,1,2$) of the resulting compact $4$-manifold with three $4$-dimensional handlebodies (of genus $k_{i}$, for $i=0,1,2$).
@@ 2.3 Gems
  • if two vertices of $\Gamma$ are $c$-adjacent ($c\in\Delta_{n}$), glue the corresponding $n$-simplices along their $(n-1)$-dimensional faces opposite to the $c$-labelled vertices, so that equally labelled vertices are identified.
@@
By construction, $K(\Gamma)$ is endowed with a vertex-labeling by $\Delta_{n}$ that is injective on each simplex. Moreover, $\Gamma$ turns out to be the 1-skeleton of the dual complex of $K(\Gamma)$. The duality establishes a bijection between the $\{c_{1},\dots,c_{h}\}$-residues of $\Gamma$ and the $(n-h)$-simplices of $K(\Gamma)$ whose vertices are labelled by $\Delta_{n}-\{c_{1},\dots,c_{h}\}$. In particular, given an $(n+1)$-colored graph $\Gamma$, each connected component of $\Gamma_{\hat{c}}$ ($c\in\Delta_{n}$) is an $n$-colored graph representing the link of a $c$-labelled vertex of $K(\Gamma)$ in the first barycentric subdivision of $K(\Gamma).$
@@
  • $|K(\Gamma)|$ is a closed $n$-manifold iff, for each color $c\in\Delta_{n}$, all $\hat{c}$-residues of $\Gamma$ represent the $(n-1)$-sphere;
  • $|K(\Gamma)|$ is a singular [^3] A singular (PL) $n$-manifold is a closed connected $n$-dimensional polyhedron admitting a simplicial triangulation where the links of vertices are closed connected $(n-1)$-manifolds, while the links of $h$-simplices, for $h\geq 1$, are $(n-h-1)$-spheres. Vertices whose links are not $(n-1)$-spheres are called singular. The notion extends also to polyhedra associated to colored graphs, provided that links are considered in the first barycentric subdivision of $K(\Gamma)$. $n$-manifold iff, for each color $c\in\Delta_{n}$, all $\hat{c}$-residues of $\Gamma$ represent closed connected $(n-1)$-manifolds.
@@
  • $\Gamma$ may be assumed to have color $n$ as its unique possible singular color, and exactly one $\hat{n}$-residue (in this case we will say that $\Gamma$ belongs to the class $G^{(n)}_{s}$);
@@
Throughout the paper, for sake of notational simplicity, if $i\in\Delta_{n}$ is such that $g_{\hat{i}}=1$, we will write $\rho_{\varepsilon}(\Gamma_{\hat{\imath}})$ (resp. $F_{\varepsilon}(\Gamma_{\hat{\imath}})$) instead of $\rho_{\varepsilon^{\prime}}(\Gamma_{\hat{\imath}})$ (resp. $F_{\varepsilon^{\prime}}(\Gamma_{\hat{\imath}})$), $\varepsilon^{\prime}$ being the permutation induced by $\varepsilon$ on $\hat{\imath}$.
@@
For each compact $n$-manifold $M$, its gem-complexity is the non-negative integer $k(M)=p-1$, where $2p$ is the minimum order of an $(n+1)$-colored graph representing $M$.
@@ 3.1 From Kirby diagrams to gems
This move, done on a quadricolor of $\Lambda(L,d)$ corresponding to a given framed component $L_{i}$, produces a new 3-colored graph representing the 3-manifold obtained by Dehn surgery on the link obtained from $L$ by adding the complementary knot of $L_{i}$ (i.e. a 00-framed trivial knot linking the component $L_{i}$ geometrically once), that is, essentially, by cancelling that framed component [^6] Quadricolors in 4-colored graphs were originally introduced by Lins ([26]). The smoothing of a quadricolor in a gem of a closed 3-manifold can also be interpreted as the substitution, in the dual triangulation, of a solid torus with another solid torus having the same boundary. Hence, it is equivalent to performing a Dehn surgery. In the specific case of $\Lambda(L,d)$, this surgery can be identified as the one along the complementary knot of the component naturally associated with the quadricolor (see [8, Proposition 9(i)] for details)..
@@
First, let us note that, as visualized in Figure 11, the handle-decomposition encoded by $(L,d)$ can be realized by considering a collar of #$\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})$, where $m\geq 0$ is the number of dotted components of $(L,d)$, and attaching the 00- and $1$-handles on one of its boundary component, the $2$-handles on the other one. This can be done via gems in the following way.
  • Starting from $\Lambda(L,d)$, let us perform the smoothing of all quadricolors of the framed components of $(L,d)$; by what we have just observed, the result is a 4-colored graph $\tilde{\Lambda}$ representing #$\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2}).$ By doubling the 1-colored edges of $\tilde{\Lambda}$ by 4-colored ones, a 5-colored graph $\tilde{\Gamma}$ is obtained which can be easily seen to represent #$\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})\times I$, whose two boundary components are represented by the 4-colored graphs $\tilde{\Gamma}_{\hat{1}}$ and $\tilde{\Gamma}_{\hat{4}}$ respectively (see the central part of Figure 11).
  • A sequence of $\rho_{2}$-pairs of color 4 (see Section 2.3) can be detected in $\tilde{\Gamma}$, whose switchings are proved to preserve the represented manifold (topologically each switching simply produces a re-triangulation of the manifold).
  • As a consequence of the $\rho_{2}$-pairs switchings, in each subgraph corresponding to a dotted component $L_{i}$ ($i\in\{1,\ldots,m\}$) a $\rho_{3}$-pair of color 4 appears, whose endpoints are exactly $(v_{i},w_{i})$ and $(v_{i}^{\prime},w_{i}^{\prime})$. As pointed out in [8, Proposition 11], the switching of these $\rho_{3}$-pairs, realized by connecting $v_{i}$ with $v_{i}^{\prime}$ and $w_{i}$ with $w_{i}^{\prime}$ by 4-colored edges for each $i\in\{1,\ldots,m\}$, has the topological effect of an identification between the boundary of a genus $m$ 4-dimensional handlebody (that is the union of 0- and 1-handles in the handle-decomposition encoded by $(L,d)$) and the component of the boundary of #$\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})\times I$ represented by $\tilde{\Gamma}_{\hat{1}}$ (the “red” boundary in Figure 11) after the $\rho_{2}$-pairs switchings.
@@
Note that, if $r=0$, then $\Gamma(L,d)$ is also a gem of $\bar{M}$ (identified with $M$ by capping off the spherical boundary). On the other hand, if $r>0$, a gem of $\bar{M}$ can be directly obtained from $\Gamma(L,d)$ in some particular cases, namely, for example, if $\Gamma(L,d)$ contains $r$$\rho_{3}$-pairs whose switchings realize the attachment of the 3-handles.
@@
Proof. With regard to statement (i), let us first suppose $r=1$ and let $\Gamma^{\prime}$ be the 5-colored graph obtained by switching the $\rho_{3}$-pair of color $i\in\Delta_{3}$ in $\Gamma$. It is not difficult to see that $\Gamma^{\prime}$ represents a closed $4$-manifold: in fact, $\Gamma^{\prime}_{\hat{4}}$ represents the 3-sphere by hypothesis, while the only ${\hat{c}}$-residues that have been modified ($c\in\Delta_{3}$) still represent $\mathbb{S}^{3}$ since they are obtained by the corresponding ${\hat{c}}$-residues of $\Gamma$ by switching a $\rho_{2}$-pair (see Section 2.3).
@@
In order to recover $\Gamma^{\prime}$, it is now necessary to delete the subgraph formed by $x,y$ and paste the resulting hanging edges. However, this subgraph is a so-called combinatorial handle in $\tilde{\Gamma}^{\prime}$, since $x$ and $y$ share the same $\{i,4\}$-colored cycle. As a consequence of [21, Theorem 7], $\Gamma^{\prime}$ turns out to be a gem of a closed $4$-manifold $N$ such that $\bar{M}\cong|K(\tilde{\Gamma}^{\prime})|\cong(\mathbb{S}^{1}\times\mathbb{S}^{3})\#N.$
More generally, if $1\leq r^{\prime}\leq r$, we insert $r^{\prime}$ 3-dipoles in $\Gamma$ in order to obtain a new gem $\tilde{\Gamma}$ of $M$ containing $r$ $\rho_{3}$-pairs, then by switching them and deleting the $r^{\prime}$ resulting combinatorial handles, a 5-colored graph is obtained which, by the above arguments, is proved to represent a closed $4$-manifold $N$ such that, as claimed, $\bar{M}\cong\#_{r^{\prime}}(\mathbb{S}^{1}\times\mathbb{S}^{3})\#N.$
@@
Figure 13: gems of $M=\mathbb{S}^{2}\times\mathbb{D}^{2}$ (left) and $\bar{M}=\mathbb{S}^{4}$ (right), arising from the 00-framed trivial knot
@@ 3.2 From gems to trisections
As a consequence, they completed the computation of the trisection genus for all the so-called “standard” simply-connected 4-manifolds, i.e. for the connected sums of $\mathbb{CP}^{2}$ possibly with reversed orientation -, $\mathbb{S}^{2}\times\mathbb{S}^{2}$ and the $K3$-surface; and for all such manifolds the trisection genus turns out to coincide with the second Betti number.
@@ 3.2.1 Gem-induced trisections
Moreover, if $H_{12}=H_{1}\cap H_{2}$ is a 3-dimensional handlebody, too, then all the above handlebodies, as well as the surface $\Sigma$, are orientable or not according to the orientability of $M$; in the first case $\Sigma$ has genus $\rho_{\varepsilon}(\Gamma_{\hat{4}})$, while in the second one it has genus $2\rho_{\varepsilon}(\Gamma_{\hat{4}}).$
@@
If the triangulation associated to a colored graph is considered, where all 4-simplices have one vertex of each color from 00 to $4$, a tricoloring is easily obtained by giving color red to both the $\varepsilon_{0}$- and $\varepsilon_{2}$-labelled vertices, color green to both the $\varepsilon_{1}$- and $\varepsilon_{3}$-labelled vertices, and color blue to the vertex labelled by $\varepsilon_{4}=4$ (which can be actually assumed to be unique, since the boundary is either empty [^9] Obviously, in the closed case, the role of color $4$ can be played by any color $c$ such that $\Gamma_{\hat{c}}$ is connected. In this paper, we always use $4$ for simplicity, both in the closed and boundary case. or connected: see Theorem 2).
@@
On the other hand, when the boundary is non-empty, the obtained decomposition of $M$ is different from that introduced in [18], which is commonly intended as a trisection in the boundary case: in fact, it doesn’t involve compression bodies and open book decompositions, but only handlebodies and Heegaard splittings, according to a suggestion by Rubinstein and Tillmann in [33, Section 4.5].
@@
there exists $\text{there exists }i\in\Delta_{3}\text{ such that}&\ \text{all 4-colored edges of the }\\ \{4,i\}\text{-colored cycle}\ \text{containing }e_{j}&\ \text{belong to the set }\{e_{1},\ldots,e_{j}\},\end{aligned}\hskip 15.0pt(*)$
@@
The positive aspect of the above result is that the involved class possibly comprehends all simply-connected closed 4-manifolds, according to the famous Kirby problem n. 50 (see [27, Section 7.3]).

### v3

@@ 2 Preliminaries about the involved representations
Obviously ${\partial{\mathbb{Y}}}_{m}^{(\sim)}=\#_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ and, whenever orientability is assumed, the $n$-sphere (resp. the $4$-disk) will be included in the notation $_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{n-1})=\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{n-1})$ (resp. ${\mathbb{Y}}_{m}^{(\sim)}={\mathbb{Y}}_{m}$) as the case $m=0.$
@@ 2.1 Kirby diagrams of orientable $4$-manifolds
In fact, since the existence of exactly one 0-handle can always be assumed, then 1- and 2-handles can be visualized, by positioning on the boundary of the 0-handle, as follows:
@@
In [33], Kirby diagrams representing closed orientable $4$-manifolds (i.e. representing compact $4$-manifolds with boundary homeomorphic to $_{r}(\mathbb{S}^{1}\times\mathbb{S}^{2}),$ with $r\geq 0$) are also called (4-dimensional) Heegaard diagrams.
@@ 2.2 Trisections of closed $4$-manifolds
A $(g;k_{0};k_{1};k_{2})$-trisection diagram is a 4-tuple $(\Sigma;\alpha,\beta,\gamma)$ such that the triples $(\Sigma;\alpha,\beta)$, $(\Sigma;\alpha,\gamma)$ and $(\Sigma;\beta,\gamma)$ are genus $g$ Heegaard diagrams for $_{k_{0}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$, $_{k_{1}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ and $_{k_{2}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ respectively. [^2] Obviously, all $3$- and $4$-dimensional manifolds involved in Definition 2, as well as in the procedure to reconstruct the represented closed $4$-manifold, are orientable or not according to $\Sigma$.
@@
    fill the three boundary components (isomorphic to $_{k_{i}}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$, for $i=0,1,2$) of the resulting compact $4$-manifold with three $4$-dimensional handlebodies (of genus $k_{i}$, for $i=0,1,2$).
@@ 2.3 Gems
  • if two vertices of $\Gamma$ are $c$-adjacent ($c\in\Delta_{n}$), glue the corresponding $n$-simplices along their (n1)(n-1)-dimensional faces opposite to the $c$-labelled vertices, so that equally labelled vertices are identified.
@@
By construction, $K(\Gamma)$ is endowed with a vertex-labeling by $\Delta_{n}$ that is injective on each simplex. Moreover, $\Gamma$ turns out to be the 1-skeleton of the dual complex of $K(\Gamma)$. The duality establishes a bijection between the $\{c_{1},\dots,c_{h}\}$-residues of $\Gamma$ and the (nh)(n-h)-simplices of $K(\Gamma)$ whose vertices are labelled by $\Delta_{n}-\{c_{1},\dots,c_{h}\}$. In particular, given an $(n+1)$-colored graph $\Gamma$, each connected component of $\Gamma_{\hat{c}}$ ($c\in\Delta_{n}$) is an $n$-colored graph representing the link of a $c$-labelled vertex of $K(\Gamma)$ in the first barycentric subdivision of $K(\Gamma).$
@@
  • $|K(\Gamma)|$ is a closed $n$-manifold iff, for each color $c\in\Delta_{n}$, all $\hat{c}-$residues of $\Gamma$ represent the (n1)(n-1)-sphere;
  • $|K(\Gamma)|$ is a singular [^3] A singular (PL) $n$-manifold is a closed connected $n$-dimensional polyhedron admitting a simplicial triangulation where the links of vertices are closed connected (n1)(n-1)-manifolds, while the links of $h$-simplices, for $h\geq 1$, are (nh1)(n-h-1)-spheres. Vertices whose links are not (n1)(n-1)-spheres are called singular. The notion extends also to polyhedra associated to colored graphs, provided that links are considered in the first barycentric subdivision of $K(\Gamma)$. $n$-manifold iff, for each color $c\in\Delta_{n}$, all $\hat{c}-$residues of $\Gamma$ represent closed connected (n1)(n-1)-manifolds.
@@
[…]
@@
[…]
@@
For each compact $n$-manifold $M$, its gem-complexity is the non-negative integer k(M)=p1k(M)=p-1, where $2p$ is the minimum order of an $(n+1)$-colored graph representing $M$.
@@ 3.1 From Kirby diagrams to gems
This move, done on a quadricolor of $\Lambda(L,d)$ corresponding to a given framed component $L_{i}$, produces a new 4-colored graph representing the 3-manifold obtained by Dehn surgery on the link obtained from $L$ by adding the complementary knot of $L_{i}$ (i.e. a 0-framed trivial knot linking the component $L_{i}$ geometrically once), that is, essentially, by cancelling that framed component [^6] Quadricolors in 4-colored graphs were originally introduced by Lins ([27]). The smoothing of a quadricolor in a gem of a closed 3-manifold can also be interpreted as the substitution, in the dual triangulation, of a solid torus with another solid torus having the same boundary. Hence, it is equivalent to performing a Dehn surgery. In the specific case of $\Lambda(L,d)$, this surgery can be identified as the one along the complementary knot of the component naturally associated with the quadricolor (see [8, Proposition 9(i)] for details)..
@@
First, let us note that, as visualized in Figure 11, the handle-decomposition encoded by $(L,d)$ can be realized by considering a collar of $_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})$, where $m\geq 0$ is the number of dotted components of $(L,d)$, and attaching the 0- and $1$-handles on one of its boundary component, the $2$-handles on the other one. This can be done via gems in the following way.
  • Starting from $\Lambda(L,d)$, let us perform the smoothing of all quadricolors of the framed components of $(L,d)$; by what we have just observed, the result is a 4-colored graph $\tilde{\Lambda}$ representing $_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2}).$ By doubling the 1-colored edges of $\tilde{\Lambda}$ by 4-colored ones, a 5-colored graph $\tilde{\Gamma}$ is obtained which can be easily seen to represent $_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})\times I$, whose two boundary components are represented by the 4-colored graphs $\tilde{\Gamma}_{\hat{1}}$ and $\tilde{\Gamma}_{\hat{4}}$ respectively (see the central part of Figure 11).
  • A sequence of $\rho_{2}$-pairs of color 4 (see Section 2.3) can be detected in $\tilde{\Gamma}$, whose switchings are proved to preserve the represented manifold. In fact, the resulting 5-colored graph can also be obtained by suturing a sequence of wounds in $\tilde{\Lambda}$, thought of as a 5-coloured graph with boundary, and subsequently by capping it off with respect to color 1 (see [21] and [9] for a detailed description of the cited moves). The fact that these operations do not change the PL type of the represented 4-manifold (they simply produce a re-triangulation of it) is mainly due to the existence in $\tilde{\Lambda}$ of a standard sequence of dipole eliminations corresponding to the highlighted segments of arcs (see [8, Remark 5]).
  • As a consequence of the $\rho_{2}$-pairs switchings, in each subgraph corresponding to a dotted component $L_{i}$ ($i\in\{1,\ldots,m\}$) a $\rho_{3}$-pair of color 4 appears, whose endpoints are exactly $(v_{i},w_{i})$ and $(v_{i}^{\prime},w_{i}^{\prime})$. As pointed out in [8, Proposition 11], the switching of these $\rho_{3}$-pairs, realized by connecting $v_{i}$ with $v_{i}^{\prime}$ and $w_{i}$ with $w_{i}^{\prime}$ by 4-colored edges for each $i\in\{1,\ldots,m\}$, has the topological effect of an identification between the boundary of a genus $m$ 4-dimensional handlebody (that is the union of 0- and 1-handles in the handle-decomposition encoded by $(L,d)$) and the component of the boundary of $_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})\times I$ represented by $\tilde{\Gamma}_{\hat{1}}$ (the “red” boundary in Figure 11) after the $\rho_{2}$-pairs switchings.
@@
[…]
@@
Proof. With regard to statement (i), let us first suppose $r=1$ and let $\Gamma^{\prime}$ be the 5-colored graph obtained by switching the $\rho_{3}$-pair of color $i\in\Delta_{3}$ in $\Gamma$. It is not difficult to see that $\Gamma^{\prime}$ represents a closed $4$-manifold: in fact, $\Gamma^{\prime}_{\hat{4}}$ represents the 3-sphere by hypothesis, while the only $^{\hat{c}}$-residues that have been modified ($c\in\Delta_{3}$) still represent $\mathbb{S}^{3}$ since they are obtained by the corresponding $^{\hat{c}}$-residues of $\Gamma$ by switching a $\rho_{2}$-pair (see Section 2.3).
@@
In order to recover $\Gamma^{\prime}$, it is now necessary to delete the subgraph formed by $x,y$ and paste the resulting hanging edges. However, this subgraph is a so-called combinatorial handle in $\tilde{\Gamma}^{\prime}$, since $x$ and $y$ share the same $\{i,4\}$-colored cycle. As a consequence of [22, Theorem 7], $\Gamma^{\prime}$ turns out to be a gem of a closed $4$-manifold NN such that $\bar{M}\cong|K(\tilde{\Gamma}^{\prime})|\cong(\mathbb{S}^{1}\times\mathbb{S}^{3})\#N.$
More generally, if $1\leq r^{\prime}\leq r$, we insert $r^{\prime}$ 3-dipoles in $\Gamma$ in order to obtain a new gem $\tilde{\Gamma}$ of $M$ containing $r$ $\rho_{3}$-pairs, then by switching them and deleting the $r^{\prime}$ resulting combinatorial handles, a 5-colored graph is obtained which, by the above arguments, is proved to represent a closed $4$-manifold NN such that, as claimed, $\bar{M}\cong\#_{r^{\prime}}(\mathbb{S}^{1}\times\mathbb{S}^{3})\#N.$
@@
Figure 13: gems of $M=\mathbb{S}^{2}\times\mathbb{D}^{2}$ (left) and $\bar{M}=\mathbb{S}^{4}$ (right), arising from the 0-framed trivial knot
@@ 3.2 From gems to trisections
As a consequence, they completed the computation of the trisection genus for all the so-called “standard” simply-connected 4-manifolds, i.e. for the connected sums of $\mathbb{CP}^{2}$ - possibly with reversed orientation -, $\mathbb{S}^{2}\times\mathbb{S}^{2}$ and the $K3$-surface; and for all such manifolds the trisection genus turns out to coincide with the second Betti number.
@@ 3.2.1 Gem-induced trisections
[…]
@@
If the triangulation associated to a colored graph is considered, where all 4-simplices have one vertex of each color from 0 to $4$, a tricoloring is easily obtained by giving color red to both the $\varepsilon_{0}$- and $\varepsilon_{2}$-labelled vertices, color green to both the $\varepsilon_{1}$- and $\varepsilon_{3}$-labelled vertices, and color blue to the vertex labelled by $\varepsilon_{4}=4$ (which can be actually assumed to be unique, since the boundary is either empty [^9] Obviously, in the closed case, the role of color $4$ can be played by any color $c$ such that $\Gamma_{\hat{c}}$ is connected. In this paper, we always use $4$ for simplicity, both in the closed and boundary case. or connected: see Theorem 2).
@@
On the other hand, when the boundary is non-empty, the obtained decomposition of $M$ is different from that introduced in [18], which is commonly intended as a trisection in the boundary case: in fact, it doesn’t involve compression bodies and open book decompositions, but only handlebodies and Heegaard splittings, according to a suggestion by Rubinstein and Tillmann in [34, Section 4.5].
@@
[…]
@@
The positive aspect of the above result is that the involved class possibly comprehends all simply-connected closed 4-manifolds, according to the famous Kirby problem n. 50 (see [28, Section 7.3]).